HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

HIGH-RESOLUTION SPECTROSCOPY OF THE GROUND AND LOW-LYING EXCITED STATES OF MgNe^{+} AND MgXe^{+}

\underline{C. KREIS} , J. R. SCHMITZ and F. MERKT, Department of Chemistry and Applied Biosciences, ETH Zürich, Wolfgang-Pauli-Strasse 10, 8093 Zürich, Switzerland

Diatomic molecules RgM consisting of a rare-gas atom Rg and an alkaline-earth-metal atom M and their singly and doubly-charged cations RgM^{+} and RgM^{2+} have unusual chemical properties related to the low first and second ionization energies of M and the high ionization energy of Rg. The second ionization energy of Mg is lower than the first ionization energy of Ar. Consequently, MgAr^{2+} is thermodynamically stable and Rydberg series of MgAr^{+} can be observed that converge on the X^{2+} ^{1}Σ^{+} ground state of MgAr^{2+} [1]. Similar behavior is expected in MgNe because of the even higher first ionization energy of Ne. In MgKr and MgXe, however, the second ionization energy of Mg is higher than the first ionization energy of the rare gas, and MgKr^{2+} and MgXe^{2+} are predicted to be metastable. In this contribution, we focus on the two extreme cases MgNe and MgXe. We present the results of spectroscopic investigations of MgNe^{+} and MgXe^{+} in their ground and low-lying electronically excited states that extend earlier studies of these cations [2,3].

PFI-ZEKE photoelectron spectra of the X ^{2}Σ^{+} ground state were recorded from the a ^{3}Π_{0} metastable state. Spectra of the lowest vibrational levels of MgNe^{+} and MgXe^{+} enabled the determination of the adiabatic ionization energy of metastable MgNe and MgXe. With a resonant two-photon (1 + 1′) excitation scheme, the PFI-ZEKE photoelectron spectrum of the X ^{2}Σ^{+} state starting from the X ^{1}Σ^{+} ground state of MgNe was recorded as well, giving access to the adiabatic ionization energy of the neutral ground state and the lowest singlet-triplet interval in the neutral molecule. Using isolated-core multiphoton Rydberg dissociation spectroscopy [4], transitions to electronically excited states of MgNe^{+} and MgXe^{+} were observed that are associated with the Rg + Mg^{+}(3p) dissociation limit. These states are the lowest members of Rydberg series converging on the ground state of MgNe^{2+} and MgXe^{2+}. These studies represent the first steps towards studying the doubly charged cations MgNe^{2+} and MgXe^{2+}.

[1] D. Wehrli, M. Génévriez, and F. Merkt, Phys. Chem. Chem. Phys., 23, 10978, (2021) [2] J. E. Reddic and M. A. Duncan, J.Chem. Phys. 110, 9948, (1999)

[3] J. S. Pilgrim, C. S. Yeh, K. R. Berry, and M. A. Duncan, J. Chem. Phys., 100, 7945, (1994) [4] M. Génévriez, D. Wehrli, and F. Merkt, Mol. Phys. 118, e1703051 (2019)

p-number: p007  Submitted on Tue, 23 May 2023 18:04:57 +0200